the perimeter of a rectangle is 40cm. If the length is 4cm greater than the width, what are the dimensions of the rectangle?
Answer by mananth(16946)

let thwe width be x
the length will be x+4
..
Perimeter = 2*(L+W)
= 2(x+x+4)
=2(2x+4)= 40
4x+8=40
add-8
4x+8-8=40-8
4x=32
divide by 4
4x/4=32/4
x=8 cm the width
length = x+4 = 12 cm